OpenStudy (anonymous):

4. Transform this equation to slope-intercept form: 6x + 4y = 8

OpenStudy (anonymous):

y = -3x/2 + 2

OpenStudy (anonymous):

You just want to solve for y 6x+4y = 8; subtract 6x from each side 6x-6x+4y = 8-6x; simplify 4y = 8-6x; divide everything by 4 4y/4 = 8/4 -6x/4; simplify y = 2-6x/4; reduce and rearrange y = -3x/2+2 <-------Answer

OpenStudy (anonymous):

could someone graph it? http://www.wolframalpha.com

OpenStudy (ash2326):

Standard equation in slope intercept form \[y=mx+c\] we have \[6x+4y=8\] subtract both side by 6x, we get \[4y=8-6x\] divide both sides by 4, we get \[y=-\frac{6}{4}x+2\] simplifying we get \[y=-\frac{3}{2} x+2\] slope= -3/2 y-intercept=2

OpenStudy (anonymous):

okay that make sense @ash2326

OpenStudy (ash2326):

Graphing is easy, it's negative slope and y intercept =2 let's find value of x when y=0 we get \[0=-3/2 x+2\] \[x=4/3\] so when x=0 y=2 and when y=0 x=4/3